# Cascading from the winds to the disk: the universality of supernovae-driven turbulence in different galactic interstellar media

### v2

@@ 1 Introduction
This study is organized as follows. In Section 2 we describe the key aspects of the multiphase gravito-hydrodynamical fluid models we use to simulate a section of a galactic disk, including the key parameters controlling the differences between each of the models (SNe-driving rate, gravitational potential and SNe seeding prescription). In Section 3 we show the key ISM statistics coming from each of the simulations, including the vertical profiles of density and turbulence, the ISM phase diagrams, and the mass density distribution functions. Next, in Section 4 we show and discuss the Fourier power spectrum for a variety of quantities, including the velocity, compressible and incompressible velocity modes, the kinetic energy and the sound speed. Finally, in Section 5 we conclude, summarize and itemize the key results in this study.
@@
Note. — Column (1): Numerical model label, where MW means Milky Way analogue, SB means starburst analogue, SC refers to the self-consistent SNe seeding prescription, Section 2.2.2, and FX refers to the fixed seeding prescription, Section 2.2.2. Column (2): Length of the cubic simulation domain. Column (3): Stellar disk component of $\phi(z)$ Equation 14. Column (4): Dark matter halo component of $\phi(z)$. Column (5): Stellar disk scale height. Column (6): Initial density. Column (7): Initial temperature. Column (8): Global, volumetric supernova driving rate. Column (9): $z$ threshold for supernova-driving boundary (FX models only; Equation 7).
2 Numerical Methods & Models
In this section, we summarize the fluid model and simulation parameters used to investigate the nature of interstellar turbulence driven by different supernova seeding schemes within varying galactic models. Our study focuses on two galaxy models, which are detailed in Section 2.1.2. One model is designed to resemble the present-day Milky Way (MW), while the other represents a starburst galaxy (SB), chosen to approximate the MW and ULTRA-MW models discussed in Martizzi et al. (2016).
2.1 Supernovae-driven, gravito-hydrodynamical fluid model
@@
where $\otimes$ is the tensor product, such that $\bm{u}\otimes\bm{u}=u_{i}u_{j}$. $\bm{u}$ is the gas velocity and $\rho$ is the mass density. $\bm{\nabla}\phi$ describes the static gravitational potential, the details of this potential are described in Section 2.1.1. $P$ is the scalar pressure, and $\mathbb{I}$ is the unit tensor, $\delta_{ij}$. $\rho e$ is the total energy composed of both the gas kinetic energy, $\rho u^{2}/2$, and the internal energy, $\epsilon$, which is related to $P$ via, $P=(\gamma-1)\rho\epsilon$, where we use $\gamma=5/3$ as the adiabatic index for a monoatomic gas. $\dot{n}_{\rm SNe}$ is the volumetric rate of SNe explosions, which changes for the two seeding schemes (details in Section 2.2), $\bm{p}_{\rm SNe}$, $M_{\rm ej}$ and $E_{\rm{th,SNe}}$ are the radial momentum, ejecta mass and thermal energy of each SNe. $M_{\rm swept}$ is the mass of ISM material swept up by the SNe shock wave. $Z$ is the metallicity of the ambient medium, which we set to $Z=Z_{\odot}$ for simplicity. Further details of the SNe driving are discussed in Section 2.2. $\Lambda$ is the net-cooling function, encompassing both heating and cooling terms, summarized in Section 2.1.3 (see Beattie et al. 2025c for details). $n_{H}$ is the number density of hydrogen. Our goal with this setup is to perform controlled studies of the coupling between SNe and turbulence in the ISM, not to implement the most complex and complete physics possible. For a detailed discussion of the limitations our numerical model we refer to §7.2 in Beattie et al. (2025c).
2.1.1 Gravitational potential
@@
2.1.2 Initial & boundary conditions
@@
Note. — All reported values in this table are computed and averaged over the statistically stationary state of the SNe-driven turbulence (Section 2.3), unless otherwise stated. The $1\sigma$ stated in the table encapsulate the physical variation of the quantities in the stationary state. Column (1): the model label, with an additional suffix for the linear grid resolution of the simulation (e.g., MWFX512 means Milky Way analogue, fixed SNe seeding prescription, discretized on a $512^{3}$ domain). Column (2): the gaseous scale height fit via a pressure-balanced isothermal atmospheric model based on our $\phi$, Equation 17. Column (3): the turbulent turnover time, Equation 12, on $\ell_{0}$. Column (4): The turbulent Mach number, Equation 13. Column (5): the midplane mass density. Column (6): the midplane pressure. Column (7): the midplane temperature. Column (8): the isothermal dispersion supporting the disk component of each of the galaxy models, fit using our isothermal profile model Equation 17. It has contributions from both the turbulence, $P_{\rm turb}\approx 1/2\bm{\nabla}u^{2}$ and the underlying thermal fluctuations.
2.1.3 Heating and cooling
We implement the heating and cooling model comprehensively described in Beattie et al. (2025c), which is the same for all simulations. To summarize, based on Theuns et al. (1998) and Sutherland & Dopita (1993), we solve a time-dependent HI, HII, HeI, HeII, HeIII and free electron chemical network at each integration step, in each grid cell based on density and temperature, using microphysical prescriptions that include photoheating and a range of cooling processes: collisional ionization, recombination, dielectronic recombination, collisional excitation, Bremsstrahlung and metal line cooling at solar metallicity. At $10^{4}\,\rm{K}\lesssim$ $T$ $\lesssim 10^{8}\,\rm{K}$, cooling is dominated by metal line emission from SNe, collisional excitation and ionization, and at $T\gtrsim 10^{7}\,\rm{K}$ bremsstrahlung dominates (Karpov et al., 2020). Since we focus primarily on the large-scales due to our resolution limits, to save the computational cost, most cooling terms truncate at $T\approx 10^{4}\,\rm{K}$, so there is no stable cold $T\approx 100\,\rm{K}$ phase. The cooling below this limit (the simulations form dilute gas with $T\lesssim 10^{2}\,\rm{K}$) is due to macroscopic processes, like adiabatic expansion of the disk (see Figure 5). At this scale, attempts to model cold phase gas (which condense in $\lesssim(0.1-10)\,\rm{pc}$ clumps (Ferrière, 2020), would be unresolved and unlikely to have a significant effect on our results. As will be seen in Section 3, our model is well-suited to capture the large-scale warm (WNM & WIM) and hot (HIM) phases of the ISM, i.e., the volume-filling properties of the disk and wind connection in the ISM.
2.2 Supernova seeding prescriptions
@@ 2.2.1 Fixed seeding (FX)
where $\dot{\Sigma}_{*}\propto\Sigma_{\rm{gas}}^{1.4}$ is the surface density of star formation (Kennicutt, 1998; Kennicutt et al., 2007). This simple seeding prescription allows SNe to explode with uniform probability inside of our steady-state gaseous disk measured in Section 3.2. Given this mechanism is simple and not linked in detail to a particular physical model, it is useful for comparison to the following more comprehensive self-consistent SNe seeding method.
@@
2.2.2 Self-consistent seeding (SC)
@@ 2.3 Statistical stationarity
where $\ell_{0}$ is the scale height of the gaseous disk in steady state, detailed in Section 3.2. Heuristically, this can be thought of as the time it takes for a turbulent fluctuation to transverse the scale height, but as demonstrated later in Section 4.4 and in Beattie et al. (2025c), it is not the outer scale of the turbulence cascade, which extends well into the winds. As discussed in Grete et al. (2025), it is not obvious that we can associate this with the time it takes the turbulence to decorrelate on this scale (Strouhal number =1=1).
@@ 3 ISM diversity
3.1 Turbulent Mach number
Along with determining the characteristic timescales in our simulations in Section 2.3, Figure 1 confirms that the volume-integral properties of the turbulence in these simulations align with approximately observations. Volume-averaged over the entire ISM, we find $\mathcal{M}_{\rm{SC}}=1.8\pm 0.1$ and $\mathcal{M}_{\rm{FX}}=1.7\pm 0.1$ for the MW model (consistent within $1\sigma$), in agreement with radio observations of the transonic WIM in the Milky Way ($\mathcal{M}\approx 2$; Gaensler et al., 2011b).
@@ 3.1 Turbulent Mach number
As noted in Kolborg et al. (2022, 2023) and above, individual SN explosions can appear as visible spikes in $\mathcal{M}$. However, we note that these fluctuations decrease drastically in intensity from $N_{\rm{grid}}=128$ to $N_{\rm{grid}}=512$. This is likely a numerical effect, due to the reduction of the volume-filling factor for the SN seeds, as injections move to progressively smaller scales with increased resolution (see Section 2.2), e.g., the SN explosions takes up a significant portion of the disk at low $N_{\rm{grid}}$, whereas the resolved WIM takes up more of the volume as $N_{\rm{grid}}$ increases. This effect is especially strong for the SB model, which we show in Figure 2 & Section 3.2 has a much thinner, more compact disk, and also has a higher global $\gamma_{\rm SNe}$ (Table 1), together leading to a well-developed wind outflow.
3.2 Scale height & galaxy profiles
The necessary condition for calculating the characteristic timescale, $t_{0}$ (Equation 12), is the definition of a characteristic length scale. We measure the scale height of the gaseous disk in steady state after evolution from initial conditions, as opposed to the aforementioned $z_{\rm{eff}}$, which is an analytical scale height derived from initial conditions to determine $z_{\rm{SNe}}$ for the FX scheme in Section 2.2.1, not including any SNe feedback. We do so by fitting an analytical model, detailed below, to the time-averaged density profile of each simulation in Figure 3.
@@ 3.2 Scale height & galaxy profiles
The velocity dispersions predicted by both SB seeding models are consistent with observational findings for the ionized gas in starburst galaxy M82, as reported by Westmoquette et al. (2009) and discussed in Section 3.1. The velocity dispersions for the MW models are $\approx 10\,\rm{km\,s}^{-1}$ in the disk, and reach $\approx 32\,\rm{km\,s}^{-1}$. This aligns with the observational findings of ionized gas in the inner $50-100\rm{pc}$ of our Galaxy when results are transformed into 3D volume-weighted values (Langer et al., 2021, $7-14\,\rm{km\,s}^{-1}$), and is also also consistent with previous expectations and simulation results (Ferrière, 2020; Gent et al., 2013). The FX runs overall have larger thermal fluctuations and a more dynamic wind, because a high fraction of SNe are exploding in low-density regions (winds and disk boundaries), efficiently heating the gas and driving outflow. This is in contrast to the SC runs, where many SNe are seeded in the dense midplane, undergoing significant radiative losses before energy can contribute efficiently to the winds. Furthermore, it should be noted that for both FX runs, there is a visible negative turning point in the slope of $\langle u^{2}\rangle^{1/2}$ and $\langle c_{s}\rangle^{1/2}$, which is $z\approx z_{\rm{SNe}}$ and SNe are no longer seeded.
As mentioned in Section 3.1, there exists a strong covariance between $\langle u^{2}\rangle^{1/2}$ and $\langle c_{s}^{2}\rangle^{1/2}$, resulting in a mostly transsonic $\mathcal{M}\approx 0.5-2$ throughout the $z$ profile. For SBFX, we overlay for comparison the Chevalier & Clegg (1985) solution for adiabatic starburst galaxy winds (turquoise dashed), which yields a scaling of $\rho\propto z^{-2}$ at $z>R$, where R is the region of mass and energy production, which we set to $Z_{\rm{SNe}}$. They assume a perfectly adiabatic, spherically symmetric flow with no external potential and constant velocity, differing from our model, which may explain the slight deviation, likely most affected by $\bm{\nabla}\phi$ and non-adiabatic processes (see Figure 6; SBFX).
3.3 Mass density distribution functions
@@
3.4 Interstellar medium phases
To better understand the relationship between our cooling function, the density structure, and the ISM phases in our simulations, as well as further compare the ISM properties of our models, we plot the time-averaged phase diagrams of $P$ and $T$ as a function of $\rho$, all normalized by their respective average values at the midplane ($z=0$; see Table 2). Isotherms are plotted at $T=10^{2}\,\rm{K}$, $10^{4}\,\rm{K}$, and $10^{6}\,\rm{K}$, and as mentioned in Section 2.1.3, the volume-filling WNM, WIM, and HIM phases are the focus of our study.
For all models, there is a clear build-up of flat probability density below the $T=10^{4}\,\rm{K}$ isotherm around the disk midplane which lies (on average) at T=3, [^0] T=3,\!000-$3,\!500\,\rm{K}$ (see Table 2). The region above $T=10^{4}\,\rm{K}$ is primarily the warm-hot wind and outflow region from SNe feedback. We have plotted an example temperature map for the SBSC model in Figure 5, masked to show material at $T>10^{4}\,\rm{K}$ (primarily winds) and $T<10^{4}\,\rm{K}$ (mostly disk), to verify the strong temperature boundaries present between the winds and disk. As mentioned in Section 2.1.3, our cooling function mostly truncates at $T=10^{4}\,\rm{K}$, but material can still cool below this from other mechanisms, most efficiently via adiabatic expansion of the galaxy disc and SNe therein (see Equation 27). In other words, nearly all the material below $T=10^{4}\,\rm{K}$ follows an adiabat, and almost all this material is in the disk. This adiabatic trajectory, $P\propto\rho^{\gamma}$ (where $\gamma=5/3$ for monoatomic gas), is most directly visible in the phase space paths from $T=10^{4}\,\rm{K}$ to $T=10^{2}\,\rm{K}$, one of which is plotted as a blue dashed line on the top-left plot. Any deviation from these adiabatic trajectories is likely due to the non-zero cooling function having small effects in the lower temperature phases, detailed in Section 2.1.3. We also see faint adiabatic cooling tracks in the WIM and HIM, due to the expansion of SNR in the diffuse winds.

### v3

@@ I Introduction
This study is organized as follows. In Section II we describe the key aspects of the multiphase gravito-hydrodynamical fluid models we use to simulate a section of a galactic disk, including the key parameters controlling the differences between each of the models (SNe-driving rate, gravitational potential and SNe seeding prescription). In Section III we show the key ISM statistics coming from each of the simulations, including the vertical profiles of density and turbulence, the ISM phase diagrams, and the mass density distribution functions. Next, in Section IV we show and discuss the Fourier power spectrum for a variety of quantities, including the velocity, compressible and incompressible velocity modes, the kinetic energy and the sound speed. Finally, in Section V we conclude, summarize and itemize the key results in this study.
@@
Note. — Column (1): Numerical model label, where MW means Milky Way analogue, SB means starburst analogue, SC refers to the self-consistent SNe seeding prescription, Section II.2.2, and FX refers to the fixed seeding prescription, Section II.2.2. Column (2): Length of the cubic simulation domain. Column (3): Stellar disk component of $\phi(z)$ Equation 14. Column (4): Dark matter halo component of $\phi(z)$. Column (5): Stellar disk scale height. Column (6): Initial density. Column (7): Initial temperature. Column (8): Global, volumetric supernova driving rate. Column (9): $z$ threshold for supernova-driving boundary (FX models only; Equation 7).
II Numerical Methods & Models
In this section, we summarize the fluid model and simulation parameters used to investigate the nature of interstellar turbulence driven by different supernova seeding schemes within varying galactic models. Our study focuses on two galaxy models, which are detailed in Section II.1.2. One model is designed to resemble the present-day Milky Way (MW), while the other represents a starburst galaxy (SB), chosen to approximate the MW and ULTRA-MW models discussed in Martizzi et al. (2016).
II.1 Supernovae-driven, gravito-hydrodynamical fluid model
@@
where $\otimes$ is the tensor product, such that $\bm{u}\otimes\bm{u}=u_{i}u_{j}$. $\bm{u}$ is the gas velocity and $\rho$ is the mass density. $\bm{\nabla}\phi$ describes the static gravitational potential, the details of this potential are described in Section II.1.1. $P$ is the scalar pressure, and $\mathbb{I}$ is the unit tensor, $\delta_{ij}$. $\rho e$ is the total energy composed of both the gas kinetic energy, $\rho u^{2}/2$, and the internal energy, $\epsilon$, which is related to $P$ via, $P=(\gamma-1)\rho\epsilon$, where we use $\gamma=5/3$ as the adiabatic index for a monoatomic gas. $\dot{n}_{\rm SNe}$ is the volumetric rate of SNe explosions, which changes for the two seeding schemes (details in Section II.2), $\bm{p}_{\rm SNe}$, $M_{\rm ej}$ and $E_{\rm{th,SNe}}$ are the radial momentum, ejecta mass and thermal energy of each SNe. $M_{\rm swept}$ is the mass of ISM material swept up by the SNe shock wave. $Z$ is the metallicity of the ambient medium, which we set to $Z=Z_{\odot}$ for simplicity. Further details of the SNe driving are discussed in Section II.2. $\Lambda$ is the net-cooling function, encompassing both heating and cooling terms, summarized in Section II.1.3 (see Beattie et al. 2025c for details). $n_{H}$ is the number density of hydrogen. Our goal with this setup is to perform controlled studies of the coupling between SNe and turbulence in the ISM, not to implement the most complex and complete physics possible. For a detailed discussion of the limitations our numerical model we refer to §7.2 in Beattie et al. (2025c).
II.1.1 Gravitational potential
@@
II.1.2 Initial & boundary conditions
@@
Note. — All reported values in this table are computed and averaged over the statistically stationary state of the SNe-driven turbulence (Section II.3), unless otherwise stated. The $1\sigma$ stated in the table encapsulate the physical variation of the quantities in the stationary state. Column (1): the model label, with an additional suffix for the linear grid resolution of the simulation (e.g., MWFX512 means Milky Way analogue, fixed SNe seeding prescription, discretized on a $512^{3}$ domain). Column (2): the gaseous scale height fit via a pressure-balanced isothermal atmospheric model based on our $\phi$, Equation 17. Column (3): the turbulent turnover time, Equation 12, on $\ell_{0}$. Column (4): The turbulent Mach number, Equation 13. Column (5): the midplane mass density. Column (6): the midplane pressure. Column (7): the midplane temperature. Column (8): the isothermal dispersion supporting the disk component of each of the galaxy models, fit using our isothermal profile model Equation 17. It has contributions from both the turbulence, $P_{\rm turb}\approx 1/2\bm{\nabla}u^{2}$ and the underlying thermal fluctuations.
II.1.3 Heating and cooling
We implement the heating and cooling model comprehensively described in Beattie et al. (2025c), which is the same for all simulations. To summarize, based on Theuns et al. (1998) and Sutherland & Dopita (1993), we solve a time-dependent HI, HII, HeI, HeII, HeIII and free electron chemical network at each integration step, in each grid cell based on density and temperature, using microphysical prescriptions that include photoheating and a range of cooling processes: collisional ionization, recombination, dielectronic recombination, collisional excitation, Bremsstrahlung and metal line cooling at solar metallicity. At $10^{4}\,\rm{K}\lesssim$ $T$ $\lesssim 10^{8}\,\rm{K}$, cooling is dominated by metal line emission from SNe, collisional excitation and ionization, and at $T\gtrsim 10^{7}\,\rm{K}$ bremsstrahlung dominates (Karpov et al., 2020). Since we focus primarily on the large-scales due to our resolution limits, to save the computational cost, most cooling terms truncate at $T\approx 10^{4}\,\rm{K}$, so there is no stable cold $T\approx 100\,\rm{K}$ phase. The cooling below this limit (the simulations form dilute gas with $T\lesssim 10^{2}\,\rm{K}$) is due to macroscopic processes, like adiabatic expansion of the disk (see Figure 5). At this scale, attempts to model cold phase gas (which condense in $\lesssim(0.1-10)\,\rm{pc}$ clumps (Ferrière, 2020), would be unresolved and unlikely to have a significant effect on our results. As will be seen in Section III, our model is well-suited to capture the large-scale warm (WNM & WIM) and hot (HIM) phases of the ISM, i.e., the volume-filling properties of the disk and wind connection in the ISM.
II.2 Supernova seeding prescriptions
@@ II.2.1 Fixed seeding (FX)
where $\dot{\Sigma}_{*}\propto\Sigma_{\rm{gas}}^{1.4}$ is the surface density of star formation (Kennicutt, 1998; Kennicutt et al., 2007). This simple seeding prescription allows SNe to explode with uniform probability inside of our steady-state gaseous disk measured in Section III.2. Given this mechanism is simple and not linked in detail to a particular physical model, it is useful for comparison to the following more comprehensive self-consistent SNe seeding method.
@@
II.2.2 Self-consistent seeding (SC)
@@ II.3 Statistical stationarity
where $\ell_{0}$ is the scale height of the gaseous disk in steady state, detailed in Section III.2. Heuristically, this can be thought of as the time it takes for a turbulent fluctuation to transverse the scale height, but as demonstrated later in Section IV.4 and in Beattie et al. (2025c), it is not the outer scale of the turbulence cascade, which extends well into the winds. As discussed in Grete et al. (2025), it is not obvious that we can associate this with the time it takes the turbulence to decorrelate on this scale (Strouhal number =1=1).
@@ III ISM diversity
III.1 Turbulent Mach number
Along with determining the characteristic timescales in our simulations in Section II.3, Figure 1 confirms that the volume-integral properties of the turbulence in these simulations align with approximately observations. Volume-averaged over the entire ISM, we find $\mathcal{M}_{\rm{SC}}=1.8\pm 0.1$ and $\mathcal{M}_{\rm{FX}}=1.7\pm 0.1$ for the MW model (consistent within $1\sigma$), in agreement with radio observations of the transonic WIM in the Milky Way ($\mathcal{M}\approx 2$; Gaensler et al., 2011b).
@@ III.1 Turbulent Mach number
As noted in Kolborg et al. (2022, 2023) and above, individual SN explosions can appear as visible spikes in $\mathcal{M}$. However, we note that these fluctuations decrease drastically in intensity from $N_{\rm{grid}}=128$ to $N_{\rm{grid}}=512$. This is likely a numerical effect, due to the reduction of the volume-filling factor for the SN seeds, as injections move to progressively smaller scales with increased resolution (see Section II.2), e.g., the SN explosions takes up a significant portion of the disk at low $N_{\rm{grid}}$, whereas the resolved WIM takes up more of the volume as $N_{\rm{grid}}$ increases. This effect is especially strong for the SB model, which we show in Figure 2 & Section III.2 has a much thinner, more compact disk, and also has a higher global $\gamma_{\rm SNe}$ (Table 1), together leading to a well-developed wind outflow.
III.2 Scale height & galaxy profiles
The necessary condition for calculating the characteristic timescale, $t_{0}$ (Equation 12), is the definition of a characteristic length scale. We measure the scale height of the gaseous disk in steady state after evolution from initial conditions, as opposed to the aforementioned $z_{\rm{eff}}$, which is an analytical scale height derived from initial conditions to determine $z_{\rm{SNe}}$ for the FX scheme in Section II.2.1, not including any SNe feedback. We do so by fitting an analytical model, detailed below, to the time-averaged density profile of each simulation in Figure 3.
@@ III.2 Scale height & galaxy profiles
The velocity dispersions predicted by both SB seeding models are consistent with observational findings for the ionized gas in starburst galaxy M82, as reported by Westmoquette et al. (2009) and discussed in Section III.1. The velocity dispersions for the MW models are $\approx 10\,\rm{km\,s}^{-1}$ in the disk, and reach $\approx 32\,\rm{km\,s}^{-1}$. This aligns with the observational findings of ionized gas in the inner $50-100\rm{pc}$ of our Galaxy when results are transformed into 3D volume-weighted values (Langer et al., 2021, $7-14\,\rm{km\,s}^{-1}$), and is also also consistent with previous expectations and simulation results (Ferrière, 2020; Gent et al., 2013). The FX runs overall have larger thermal fluctuations and a more dynamic wind, because a high fraction of SNe are exploding in low-density regions (winds and disk boundaries), efficiently heating the gas and driving outflow. This is in contrast to the SC runs, where many SNe are seeded in the dense midplane, undergoing significant radiative losses before energy can contribute efficiently to the winds. Furthermore, it should be noted that for both FX runs, there is a visible negative turning point in the slope of $\langle u^{2}\rangle^{1/2}$ and $\langle c_{s}\rangle^{1/2}$, which is $z\approx z_{\rm{SNe}}$ and SNe are no longer seeded.
As mentioned in Section III.1, there exists a strong covariance between $\langle u^{2}\rangle^{1/2}$ and $\langle c_{s}^{2}\rangle^{1/2}$, resulting in a mostly transsonic $\mathcal{M}\approx 0.5-2$ throughout the $z$ profile. For SBFX, we overlay for comparison the Chevalier & Clegg (1985) solution for adiabatic starburst galaxy winds (turquoise dashed), which yields a scaling of $\rho\propto z^{-2}$ at $z>R$, where R is the region of mass and energy production, which we set to $Z_{\rm{SNe}}$. They assume a perfectly adiabatic, spherically symmetric flow with no external potential and constant velocity, differing from our model, which may explain the slight deviation, likely most affected by $\bm{\nabla}\phi$ and non-adiabatic processes (see Figure 6; SBFX).
III.3 Mass density distribution functions
@@
III.4 Interstellar medium phases
To better understand the relationship between our cooling function, the density structure, and the ISM phases in our simulations, as well as further compare the ISM properties of our models, we plot the time-averaged phase diagrams of $P$ and $T$ as a function of $\rho$, all normalized by their respective average values at the midplane ($z=0$; see Table 2). Isotherms are plotted at $T=10^{2}\,\rm{K}$, $10^{4}\,\rm{K}$, and $10^{6}\,\rm{K}$, and as mentioned in Section II.1.3, the volume-filling WNM, WIM, and HIM phases are the focus of our study.
For all models, there is a clear build-up of flat probability density below the $T=10^{4}\,\rm{K}$ isotherm around the disk midplane which lies (on average) at T=3, [^0] T=3,\!000-$3,\!500\,\rm{K}$ (see Table 2). The region above $T=10^{4}\,\rm{K}$ is primarily the warm-hot wind and outflow region from SNe feedback. We have plotted an example temperature map for the SBSC model in Figure 5, masked to show material at $T>10^{4}\,\rm{K}$ (primarily winds) and $T<10^{4}\,\rm{K}$ (mostly disk), to verify the strong temperature boundaries present between the winds and disk. As mentioned in Section II.1.3, our cooling function mostly truncates at $T=10^{4}\,\rm{K}$, but material can still cool below this from other mechanisms, most efficiently via adiabatic expansion of the galaxy disc and SNe therein (see Equation 27). In other words, nearly all the material below $T=10^{4}\,\rm{K}$ follows an adiabat, and almost all this material is in the disk. This adiabatic trajectory, $P\propto\rho^{\gamma}$ (where $\gamma=5/3$ for monoatomic gas), is most directly visible in the phase space paths from $T=10^{4}\,\rm{K}$ to $T=10^{2}\,\rm{K}$, one of which is plotted as a blue dashed line on the top-left plot. Any deviation from these adiabatic trajectories is likely due to the non-zero cooling function having small effects in the lower temperature phases, detailed in Section II.1.3. We also see faint adiabatic cooling tracks in the WIM and HIM, due to the expansion of SNR in the diffuse winds.